# Real-time Identity Defenses against Malicious Personalization of Diffusion Models

### v1

@@ Abstract
Personalized diffusion models, capable of synthesizing highly realistic images based on a few reference portraits, pose substantial social, ethical, and legal risks by enabling identity replication. Existing defense mechanisms rely on computationally intensive adversarial perturbations tailored to individual images, rendering them impractical for real-world deployment. This study introduces Real-time Identity Defender (RID), a neural network designed to generate adversarial perturbations through a single forward pass, bypassing the need for image-specific optimization. RID achieves unprecedented efficiency, with defense times as low as 0.12 seconds on a single GPU—4,400 times faster than leading methods—and 1.1 seconds per image on a standard Intel i9 CPU, making it suitable for edge devices such as smartphones. Despite its efficiency, RID matches state-of-the-art performance across visual and quantitative benchmarks, effectively mitigating identity replication risks. Our analysis reveals that RID’s perturbations mimic the efficacy of traditional defenses while exhibiting properties distinct from natural noise, such as Gaussian perturbations. To enhance robustness, we extend RID into an ensemble framework that integrates multiple pre-trained text-to-image diffusion models, ensuring resilience against black-box attacks and post-processing techniques, including JPEG compression and diffusion-based purification.
@@ Introduction
Large-scale diffusion models have made significant strides in generating high-fidelity images from textual descriptions, driven by their ability to generalize in a zero-shot manner—synthesizing novel combinations of concepts learned from extensive pretraining datasets [1, 2, 3, 4, 5, 6] (Fig. 1a). Building on this foundation, personalization techniques[7, 8, 9, 10, 11, 12] inspired by delta tuning [13] have enabled diffusion models to replicate unique concepts, including individual identities, from as few as 10–20 reference images while maintaining generalization capabilities (Fig. 1b). This functionality has unlocked immense creative potential and found widespread applications in design, education, and entertainment.
However, the ability to replicate individual identities raises profound ethical, legal, and societal concerns. By fine-tuning open-source models like Stable Diffusion [14, 15, 5], malicious actors can generate highly realistic but harmful content, including disinformation or fabricated media, threatening personal privacy and public trust. The accessibility of pre-trained diffusion models and the efficiency of personalization techniques exacerbate these risks, with identity replication achievable in under 15 minutes using consumer-grade hardware and publicly available tools. Given the vast number of images uploaded online daily, the potential for unauthorized identity replication and manipulation is vast and urgent.
To mitigate these threats, existing defense mechanisms [16, 17, 18, 19, 20, 21] optimize protective adversarial perturbations [22, 18] for individual images to prevent diffusion models from learning the identities depicted [^1] This is a kind of adversarial attack for good, i.e. leveraging the vulnerabilities of T2I models to protect identities.. These perturbations, imperceptible to the human eye, are crafted to exploit vulnerabilities in diffusion models’ objective functions while adhering to norm constraints (Fig. 1c). Although effective, these methods are computationally intensive and impractical for large-scale or real-time applications, as they require over 500 seconds per image to perform individual optimization.
To address these limitations, we propose the Real-time Identity Defender (RID), a neural network that generates adversarial perturbations through a single forward pass, eliminating the need for image-specific optimization. RID achieves unprecedented efficiency, processing images in as little as 0.12 seconds on a single GPU—4,400 times faster than existing approaches—and 1.1 seconds per image on a standard Intel i9 CPU, enabling deployment on edge devices such as smartphones. Our approach is powered by a novel loss function, Adversarial Score Distillation Sampling (Adv-SDS), which targets pre-trained diffusion models efficiently, and a regularization term that enhances the visual quality of defended images. RID is trained on a curated dataset of 70,000 multi-person images (Fig. 1d) and achieves comparable performance to state-of-the-art methods in both visual perception and quantitative metrics while significantly improving scalability. Further analysis reveals that RID’s perturbations emulate the effectiveness of traditional defenses while remaining distinct from natural noise, such as Gaussian perturbations.
Additionally, we extend RID with an ensemble framework that leverages multiple pre-trained diffusion models, ensuring robust defenses against black-box attacks and resilience to post-processing methods such as JPEG compression and diffusion-based purification. RID offers a scalable, practical solution to safeguarding identities in the era of generative personalization, bridging the critical gap between security and usability.
@@
Figure 1: Overview of personalized TI2 diffusion models and our real-time identity defender (RID). a Personalized T2I models learn personal identity efficiently by fine-tuning a pre-trained TI2 model (e.g., Stable Diffusion [14, 15, 5]) on a few portraits. b Personalized T2I models can generate high-fidelity images by combining the learned identity and other concepts following the text prompt. c Existing defense methods optimize an individual adversarial perturbation for each image against Personalized T2I techniques. d RID trains a defender on a face dataset of multiple persons via adversarial score distillation sampling (detailed in Sec. Method). e RID defends a new testing image by generating the corresponding perturbation through an efficient forward pass. f RID successfully prevents personalized T2I techniques from learning personal identities in terms of visual perception.
@@ Results
RID introduces a neural network capable of generating protective perturbations for images in a single forward pass, eliminating the need for computationally intensive, image-specific optimization. Below, we outline the key components of the RID framework, with additional details provided in the Methods section.
@@
RID is trained on a curated dataset of 70,000 high-resolution images from the VGG-Face2 dataset [23], selected following established filtering criteria [24]. The dataset includes diverse facial features, poses, and lighting conditions, ensuring robust generalization across varied scenarios.
The primary training objective of RID is the adversarial score distillation sampling (Adv-SDS) loss, which optimizes the network to generate perturbations that disrupt the ability of personalized diffusion models to replicate protected identities. By adversarially maximizing the diffusion loss [25], Adv-SDS ensures that defended images resist personalization while maintaining visual fidelity. This approach leverages the intrinsic properties of pre-trained diffusion models to craft effective perturbations, achieving a robust trade-off between identity protection and image quality.
Technically, Adv-SDS addresses a significant computational challenge: directly calculating the Jacobian of large diffusion models with respect to their inputs is infeasible due to their complexity. Drawing inspiration from score distillation sampling (SDS)—which minimizes diffusion loss for zero-shot 3D generation [26] rather than maximizing it for protection—we propose an efficient and theoretically grounded surrogate objective for Adv-SDS. This surrogate simplifies gradient computations while maintaining the defense’s effectiveness.
Although Adv-SDS produces strong identity protection, it occasionally introduces subtle grid-like artifacts into the defended images. To address this, we incorporate a regression loss that aligns RID’s perturbations with precomputed, image-specific perturbations from Anti-DB [19]. This regularization term enhances visual consistency and reduces artifacts without significantly increasing training complexity. However, our ablation study (Appendix xx) reveals that the regression loss alone does not generalize effectively in isolation, highlighting its role as a supplementary component.
@@
Once trained, RID generates protective perturbations for input images in a single forward pass. Each perturbation is directly added to the corresponding input image to create a defended version. This process is highly efficient, as shown in Fig. 2b.
Figure 2: RID defends the identities against malicious image generation efficiently and effectively under various metrics. a The RID-defended image closely resembles the original clean image (left column). Personalized diffusion models trained on clean images accurately retain identity across three prompts (top row), while models trained on RID-defended images produce distorted outputs with reduced identifiable features (bottom row). b RID achieves significantly faster defense speeds, with processing times of 1.1 seconds on CPU and 0.12 seconds on GPU, compared to optimization-based methods such as Anti-DB and AdvDM, which require over 500 seconds on GPU. c Quantitative evaluation across three metrics (FID, ISM, and BRISQUE) shows that RID provides comparable protection to optimization-based methods and performs significantly better than the baseline without any defense. Arrows next to each metric denote better defense performance (i.e., lower visual quality). d Although the visual patterns of RID-defended samples differ from those of existing methods, RID achieves comparable defense performance qualitatively, consistent with the quantitative results in Panel c.
@@
For the evaluation, we used the CelebA-HQ dataset [27], different from the training set, to assess the generalization capabilities of RID. A total of 15 individuals (five females and ten males) were randomly selected, with 12 images per individual. During training, personalized diffusion models were fine-tuned on clean or RID-defended images using the prompt “Photo of a [*] person”. This consistent training prompt simulates real-world adversarial scenarios, where publicly available images are used to train models for identity replication.
For inference, we generated outputs under two conditions: the same prompt used during training (“Photo of a [*] person”) to assess alignment with training conditions for quantitative evaluation, and diverse prompts such as “[*] person with a dog” to evaluate RID’s robustness in generalizing to novel scenarios. This two-fold testing approach ensures a comprehensive evaluation of RID’s defense performance across varying conditions.
@@
To assess RID’s defense capabilities, we employed several personalization techniques. For the loss function, we adopted Dreambooth [28], which preserves the generative generalization of personalized models. Two representative fine-tuning approaches were considered: the first involved training all parameters of the diffusion model (referred to as DB) [28], while the second utilized a combination of two commonly used lightweight methods—textual embedding (TI) [29] and LoRA [30] (denoted LoRA+TI). DB was chosen for its strong personalization capability, whereas LoRA+TI serves as a widely adopted lightweight alternative. Notably, LoRA+TI typically outperforms TI or LoRA alone and is much more efficient than DB, offering an excellent balance between efficiency and effectiveness. Therefore, it is used as the default method in our experiments.
@@
Quantitative evaluation focused on two key aspects: identity protection and visual fidelity, using three common metrics. Frechet Inception Distance (FID)[31] measured the similarity between the distributions of generated and reference images, with higher FID scores for RID-defended images indicating stronger defense. Identity Score Matching (ISM)[19] evaluated identity similarity using ArcFace [32] embeddings, where lower ISM scores reflect better identity protection. Finally, the Blind/Referenceless Image Spatial Quality Evaluator (BRISQUE) [33] assessed perceptual image quality, with higher scores indicating greater visual degradation. Details of the metrics are provided in the Appendix.
RID defends against malicious personalization effectively and efficiently
To evaluate the effectiveness of RID, we conducted comprehensive qualitative and quantitative analyses, as shown in Figs. 2 and 3. These evaluations demonstrate RID’s ability to safeguard identities against personalization by text-to-image diffusion models while maintaining computational efficiency and adaptability across various settings.
@@
Fig.2b compares the processing speeds of RID with representative optimization-based methods, such as Anti-DB[19] and AdvDM [17]. While existing methods require over 500 seconds per image on a GPU, RID completes the defense process in just 0.12 seconds on a GPU and 1.1 seconds on a CPU. This makes RID approximately 4,400 times faster, enabling practical deployment in real-time applications, including on edge devices like smartphones, while maintaining comparable defense performance.
@@
Fig. 2c presents a quantitative evaluation of RID’s defense effectiveness under three widely used metrics: FID, ISM, and BRISQUE, which collectively assess visual quality and defense robustness. RID achieves competitive performance across all three metrics, providing comparable protection to optimization-based methods while significantly outperforming the baseline without defense. The arrows in the panel indicate the directions of better defense performance (e.g., higher FID and BRISQUE, and lower ISM), further validating RID’s effectiveness.
@@
Fig. 2d examines the visual patterns of RID-defended images in comparison to those protected by Anti-DB and AdvDM. Although the perturbation patterns generated by RID differ from those of optimization-based methods, all approaches achieve qualitatively effective protection, as evidenced by the distorted outputs shown in Panel a. This highlights the practical efficacy of RID’s defenses, even with its generalized, image-agnostic design.
RID is robust across perturbation levels, personalization techniques, and pre-trained models
Figure 3: RID is robust across perturbation levels, pre-trained diffusion models, and personalization techniques. Under the FID, ISM, and BRISQUE metrics, RID demonstrates effective defense performance compared to the baseline on clean images, across three perturbation levels (6/255, 8/255, and 12/255), popular pre-trained diffusion models (SD v2.1, v1.5, and v1.4), and representative personalization techniques (DB and LoRA+TI). In all panels, the $x$-axis represents the perturbation scale, with 00 indicating baseline results on clean images. Arrows next to each metric indicate better defense performance (i.e., lower visual quality).
As shown in Fig. 3, we evaluate RID’s defense performance across different perturbation levels, personalization techniques, and pre-trained diffusion models. These results, measured using FID, ISM, and BRISQUE metrics, demonstrate RID’s adaptability and robustness across diverse setups.
@@
Across all columns, the $x$-axis represents perturbation scales (6/255, 8/255, and 12/255), with 0 indicating baseline results on clean images. As the perturbation scale increases, defense effectiveness improves, reflected by higher FID and BRISQUE scores and lower ISM scores. Nevertheless, RID provides effective defense even at lower perturbation levels. For example, ISM scores drop from 0.59 to approximately 0.4, while FID increases from 169.26 to over 300, indicating successful disruption of identity retention.
@@
The first column corresponds to Dreambooth (DB) [28] applied to Stable Diffusion (SD) v2.1, while the second column examines LoRA+TI [29] applied to the same model. Similar trends are observed across both techniques: ISM scores decrease from 0.61 to around 0.3, FID rises from 139.72 to over 300, and BRISQUE increases from 11.91 to 44.35 as the perturbation scale increases. These results demonstrate that RID effectively disrupts identity replication and maintains consistent performance regardless of the personalization method.
@@
The third and fourth columns evaluate RID on LoRA+TI applied to SD v1.5 and SD v1.4, respectively. Together with the results for SD v2.1, RID demonstrates consistent trends across all pre-trained diffusion models, with defense performance improving as perturbation scales increase. These results further validate RID’s generalizability and effectiveness across diverse model architectures.
@@
An analysis of the perturbations generated by RID
Figure 4: Analysis of the perturbations generated by RID. The pixel values are normalized from $[0,255]$ to $[-1,1]$. Two Gaussian noise baselines with standard deviations of the perturbation level divided by $3$ (small, following the 3-sigma rule) and $1$ (large) are used for comparison, discretized to match the perturbation levels and normalized pixel range. a. Every perturbation is flattened as a vector to count the values of all elements. The distributions of RID perturbations differ markedly from those of Gaussian noise. b. Visualization of RID perturbations as images reveals structured, semantic patterns aligned with facial features, similar to those of optimization-based methods, while Gaussian noise appears random and unstructured. Perturbations are displayed at a scale of $1$ (e.g., multiplied by a factor of $255/8$ for RID) for clarity. c. RID perturbations consistently increase diffusion losses compared to Gaussian noise across various pre-trained and personalized diffusion models, suggesting the effectiveness of learned perturbations over natural noise.
This section examines the properties of perturbations generated by RID in comparison to both Gaussian noise baselines and existing optimization-based methods. As shown in Fig.4, RID produces semantically meaningful perturbations that differ markedly from Gaussian noise and behave similarly to those of optimization-based methods, such as Anti-DB [19] and AdvDM [17].
@@
The pixel distributions of RID-generated perturbations, as shown in Fig. 4a, differ significantly from those of Gaussian noise baselines. All pixel values are normalized to [-1,1], with two Gaussian baselines included for comparison: small Gaussian noise (with a standard deviation of one-third of the perturbation level, following the 3-sigma rule) and large Gaussian noise (matching the full perturbation level). RID perturbations exhibit a heavy-tailed statistical distribution, distinct from the symmetric distribution of Gaussian noise (even with truncation). This characteristic highlights RID’s ability to learn structured perturbations that deviate from the random and unstructured patterns seen in Gaussian noise.
@@
As shown in Fig. 4b, RID-generated perturbations reveal structured and semantic patterns that align closely with facial features, similar to those produced by optimization-based methods such as Anti-DB and AdvDM. These perturbations focus on regions containing meaningful information, such as facial contours, to effectively disrupt identity replication. In contrast, Gaussian noise lacks any semantic structure, appearing random and uncorrelated with the content of the image. The semantic alignment of RID perturbations reflects the learned nature of its defense mechanism, ensuring that sensitive features are targeted for protection.

### v2

# Protecting your portraits: Real-time Identity Defense against malicious personalization of diffusion models

@@ Abstract
Personalized generative diffusion models, capable of synthesizing highly realistic images based on a few reference portraits, may pose substantial social, ethical, and legal risks via identity replication. Existing defense mechanisms rely on computationally intensive adversarial perturbations tailored to individual images, rendering them impractical for real-world deployment. This study introduces the Real-time Identity Defender (RID), a neural network designed to generate adversarial perturbations through a single forward pass, bypassing the need for image-specific optimization. RID achieves unprecedented efficiency, with defense times as low as 0.12 seconds on a single NVIDIA A100 80G GPU (4,400 times faster than leading methods) and 1.1 seconds per image on a standard Intel i9 CPU, making it suitable for edge devices such as smartphones. Despite its efficiency, RID achieves promising protection performance across visual and quantitative benchmarks, effectively mitigating identity replication risks. Our analysis reveals that RID’s perturbations mimic the efficacy of traditional defenses while exhibiting properties distinct from natural noise, such as Gaussian perturbations. To enhance robustness, we extend RID into an ensemble framework that integrates multiple pre-trained text-to-image diffusion models, ensuring resilience against black-box attacks and post-processing techniques, including image compression and purification. Our model is envisioned to play a crucial role in safeguarding portrait rights, thereby preventing illegal and unethical uses.
@@ Introduction
Large-scale diffusion models have made significant strides in generating high-fidelity images from textual descriptions, driven by their ability to generalize in a zero-shot manner—synthesizing novel combinations of concepts learned from extensive pretraining datasets [1, 2, 3, 4, 5, 6] (see Fig. 1a). Building on this foundation, personalization techniques [7, 8, 9, 10, 11, 12] inspired by delta tuning [13] have enabled diffusion models to replicate unique concepts, including individual identities, from as few as a dozen of reference images while maintaining generalization capabilities (see Fig. 1b). This functionality has unlocked immense creative potential and found widespread applications in design, education, and entertainment.
@@
Figure 1: Overview of personalized T2I diffusion models and our real-time identity defender (RID). In all panels, the flame icon Refer to caption indicates trainable parameters. a, Personalized T2I models learn personal identity efficiently by fine-tuning a pre-trained TI2 model (e.g., Stable Diffusion [14, 15, 5]) on a few portraits. b, Personalized T2I models can generate high-fidelity images by combining the learned identity and other concepts following the text prompt. c, Existing defense methods optimize an individual adversarial perturbation for each image against Personalized T2I techniques. d, RID trains a defender on a face dataset of multiple persons via adversarial score distillation sampling and regularization (detailed in the Method section). e, RID defends a new testing image by generating the corresponding perturbation through an efficient forward pass. f, RID successfully prevents personalized T2I techniques from learning personal identities in terms of visual perception. All facial images used in this figure are sourced from publicly available datasets and are permitted for academic purposes.
However, the ability to replicate individual identities raises profound ethical, legal, and societal concerns. By fine-tuning open-source models like Stable Diffusion [14, 15, 5], malicious actors can generate highly realistic but harmful content, including disinformation or fabricated media, threatening personal privacy and public trust. The accessibility of pre-trained diffusion models and the efficiency of personalization techniques amplify these risks, enabling identity replication achievable within 15 minutes using consumer-grade hardware and publicly available tools. Given the vast number of images uploaded online daily, the potential for unauthorized identity replication and manipulation is vast and urgent.
To mitigate these threats, existing defense mechanisms [16, 17, 18, 19, 20, 21] optimize protective adversarial perturbations [22, 18] for individual images, preventing diffusion models from learning the depicted identities [^1] This is a kind of adversarial attack for good, i.e., leveraging the vulnerabilities of T2I models to protect identities.. These perturbations, imperceptible to the human eye, are crafted to exploit vulnerabilities in diffusion models’ objective functions while adhering to norm constraints (see Fig. 1c). Despite effective, these methods are computationally intensive and impractical for large-scale or real-time applications, as they require over 500 seconds per image to perform individual optimization.
To address these limitations, we propose the Real-time Identity Defender (RID), a neural network that generates adversarial perturbations through a single forward pass, eliminating the need for image-specific optimization. RID achieves unprecedented efficiency, processing images in as little as 0.12 seconds on a single NVIDIA A100 80G GPU (4,400 times faster than existing approaches) and 1.1 seconds per image on a standard Intel core i9 CPU, having potential to enable deployment on edge devices such as smartphones. Our approach is powered by a novel loss function, namely, Adversarial Score Distillation Sampling (Adv-SDS), which targets pre-trained diffusion models efficiently, and a regularization term that enhances the visual quality of defended images. RID is trained on a curated dataset of 70,000 multi-person images (see Fig. 1d) and achieves promising protection performance in both visual perception and quantitative metrics while significantly improving efficiency and scalability. Our further analyses reveal that RID’s perturbations emulate the effectiveness of traditional defenses while remaining distinct from natural noise, such as Gaussian perturbations.
Additionally, we extend RID with an ensemble framework that leverages multiple pre-trained diffusion models, ensuring robust defenses against black-box attacks and resilience to post-processing methods such as JPEG compression and diffusion-based purification. As a result, RID offers a scalable, practical solution to safeguarding identities in the era of generative personalization, bridging the critical gap between security and usability.
@@ Results
Pipeline.
RID introduces a neural network capable of generating protective perturbations for images in a single forward pass, eliminating the need for computationally intensive, image-specific optimization. As illustrated in Extended Data Fig. 1, RID takes a clean image as input and outputs an adversarial noise. This noise is then added to the clean image to produce a protected image, which is subsequently fed into a pre-trained diffusion model that attempts to personalize it. Intuitively, the loss function of RID is designed to increase the difficulty of the personalization process, thereby achieving effective identity protection. After being trained on multi-person image datasets, RID will possess strong generalization capabilities, generating protective noise for new, unseen data through a single forward pass.
@@
We propose adversarial score distillation sampling (Adv-SDS) loss as the primary training objective of RID, which optimizes the network to generate perturbations that disrupt the ability of personalized diffusion models to replicate protected identities. By adversarially maximizing the diffusion loss [23], Adv-SDS ensures that defended images resist personalization while maintaining visual fidelity. This approach leverages the intrinsic properties of pre-trained diffusion models to craft effective perturbations, essentially achieving a robust trade-off between identity protection and image quality.
Technically, Adv-SDS addresses a significant computational challenge: directly calculating the Jacobian of large diffusion models with respect to their inputs is infeasible due to their complexity. Drawing inspiration from score distillation sampling (SDS), which minimizes diffusion loss for zero-shot 3D generation [24] rather than maximizing it for protection, we propose an efficient and theoretically grounded surrogate objective for Adv-SDS. This surrogate simplifies gradient computations while maintaining the defense’s effectiveness.
Although Adv-SDS produces promising identity protection, it occasionally introduces subtle grid-like artifacts into the defended images. To address this, we incorporate a regression loss that aligns RID’s perturbations with precomputed, image-specific perturbations from Anti-DB [19]. This regularization term enhances visual consistency and reduces artifacts without significantly increasing training complexity. However, our ablation study (See Extended Data Fig. 2) reveals that the regression loss alone does not protect the identity effectively in isolation, highlighting its important role as a supplementary component.
Noteworthy, RID is trained on a curated dataset of 70,000 high-resolution images from the VGG-Face2 dataset [25], selected following the established filtering criteria [26]. The dataset includes diverse facial features, poses, and lighting conditions, ensuring robust generalization across varied scenarios.
@@
Once trained, RID generates protective perturbations for input images in a single forward pass. Each perturbation is directly added to the corresponding input image to create a defended version. This process is highly efficient, as shown in Fig. 2b, achieving over three orders of magnitude speedup compared with existing methods (e.g., Anti-DB and AdvDM) under the same computational environment.
Figure 2: RID defends the identities against malicious image generation efficiently and effectively under various metrics. a, The RID-defended image closely resembles the clean image (left column). Personalized diffusion models trained on clean images accurately retain identity across three prompts (top row), while models trained on RID-defended images produce distorted outputs with reduced identifiable features (bottom row). b, RID achieves significantly faster defense speeds, with processing times of 1.1 seconds on an Intel I9 CPU and 0.12 seconds on a NVIDIA A100 80G GPU, compared to optimization-based methods such as Anti-DB and AdvDM, which require over 500 seconds on the same GPU. c, Quantitative evaluation across three metrics (FID, ISM, and BRISQUE) shows that RID provides comparable protection to optimization-based methods and performs significantly better than the baseline without any defense. Arrows next to each metric denote better defense performance (i.e., lower visual quality). d, Although the visual patterns of RID-defended samples differ from those of existing methods, RID achieves comparable defense performance qualitatively, consistent with the quantitative results depicted in c. All facial images used in this figure are sourced from publicly available datasets and are permitted for academic purposes.
@@
For the evaluation, we used the open-source CelebA-HQ dataset [27], different from the training set, to assess the generalization capabilities of RID. A total of 15 individuals (five females and ten males) were randomly selected, with 12 images per individual. During training, personalized diffusion models were fine-tuned on clean or RID-defended images using the prompt “Photo of a [*] person”. This consistent training prompt simulates real-world adversarial scenarios, where publicly available images are used to train models for identity replication.
For inference, we generated outputs under two conditions: the same prompt used during training to assess alignment with training conditions for quantitative evaluation, and diverse prompts such as “[*] person with a dog” to evaluate RID’s robustness in generalizing over novel scenarios. This two-fold testing approach ensures a comprehensive evaluation of RID’s defense performance across varying conditions.
@@
To assess RID’s defense capabilities, we employed several personalization techniques. For the loss function, we adopted Dreambooth [28], which preserves the generative generalization of personalized models. Two representative fine-tuning approaches were considered: the first involved training all parameters of the diffusion model (referred to as DB) [28], while the second utilized a combination of two commonly used lightweight methods—textual embedding (TI) [29] and LoRA [30] (denoted as LoRA+TI). DB was chosen for its strong personalization capability, whereas LoRA+TI serves as a widely adopted lightweight alternative. Notably, LoRA+TI typically outperforms TI or LoRA alone and is much more efficient than DB, offering an excellent balance between efficiency and effectiveness. Therefore, it is used as the default method throughout our experiments.
@@
The quantitative evaluation focused on two key aspects, namely, identity protection and visual fidelity, using three common metrics. Frechet Inception Distance (FID) [31] measured the similarity between the distributions of generated and reference images, with higher FID scores for RID-defended images indicating a stronger defense. Identity Score Matching (ISM) [19] evaluated identity similarity using ArcFace [32] embeddings, where lower ISM scores reflect a better identity protection. Finally, the Blind/Referenceless Image Spatial Quality Evaluator (BRISQUE) [33] assessed the perceptual image quality, with higher scores indicating a greater visual degradation. Details of the metrics are provided in the Method section.
Defending against malicious personalization
To evaluate the effectiveness of RID, we conducted comprehensive qualitative and quantitative analyses, with the results shown in Figs. 2 and 3. These evaluations demonstrate RID’s ability to safeguard identities against personalization by text-to-image diffusion models while maintaining computational efficiency and adaptability across various settings.
@@
Fig. 2b compares the processing speeds of RID with representative optimization-based methods, such as Anti-DB [19] and AdvDM [17]. While existing methods require over 500 seconds per image on a GPU, RID completes the defense process in just 0.12 seconds on a GPU and 1.1 seconds on a CPU. This makes RID approximately 4,400 times faster, possessing potential to enable practical deployment in real-time applications, such as on edge devices like smartphones, while maintaining a comparable defense performance.
@@
Fig. 2c presents a quantitative evaluation of RID’s defense effectiveness under three widely used metrics: FID, ISM, and BRISQUE, which collectively assess visual quality and defense robustness. RID achieves competitive performance across all three metrics, yielding comparable protection to optimization-based methods while significantly outperforming the baseline without defense. Although the the metrics in Fig. 2c may show slightly weaker performance of RID compared with Anti-DB and AdvDM, a comparable level of protection is achieved, as demonstrated in Fig. 2d, where visual inspection finds it difficult to distinguish the original image.
@@
Fig. 2d examines the visual patterns of RID-defended images in comparison to those protected by Anti-DB and AdvDM. Although the perturbation patterns generated by RID differ from those of optimization-based methods, all approaches achieve qualitatively effective protection, as evidenced by the distorted outputs shown in Fig. 2a. This highlights the practical efficacy of RID’s defenses, even with its generalized, image-agnostic design.
Robustness of RID
Figure 3: Robustness of RID across perturbation levels, pre-trained diffusion models, and personalization techniques. Under the FID, ISM, and BRISQUE metrics, RID demonstrates effective defense performance compared to the baseline on clean images, across three perturbation levels (6/255, 8/255, and 12/255), popular pre-trained diffusion models (SD v2.1, v1.5, and v1.4), and representative personalization techniques (DB and LoRA+TI). In all panels, the $x$-axis represents the perturbation scale, with 00 indicating baseline results on clean images. Arrows at the end of each row indicate better defense performance (i.e., lower visual quality) under the corresponding metric.
To test RID’s robustness, we evaluate RID’s defense performance across different perturbation levels, personalization techniques, and pre-trained diffusion models. The results, measured using FID, ISM, and BRISQUE metrics, demonstrate RID’s adaptability and robustness across diverse setups as shown in Fig. 3.
@@
The $x$-axes in Fig. 3 represent the perturbation scales (6/255, 8/255, and 12/255), with 0 indicating the baseline results on clean images. As the perturbation scale increases, the defense effectiveness improves, reflected by higher FID and BRISQUE scores and lower ISM scores. Nevertheless, RID provides effective defense even at lower perturbation levels. For example, ISM scores drop from 0.59 to approximately 0.4, while FID increases from 169.26 to over 300, indicating successful disruption of identity retention.
@@
We firstly test the robustness of RID over different personalization techniques (aka, DB [28] and LoRA+TI [29]), based on the pre-trained diffusion model, namely, Stable Diffusion (SD) v2.1. As shown in Fig. 3, similar trends are observed across both techniques: ISM scores decrease from 0.61 to around 0.3, FID rises from 139.72 to over 300, and BRISQUE increases from 11.91 to 44.35 as the perturbation scale increases. These results demonstrate that RID effectively disrupts identity replication and maintains consistent performance regardless of the specific personalization method.
@@
We further test the robustness of RID across different pre-trained diffusion models, including SD v1.5 and SD v1.4, by employing the LoRA+TI personalization technique. Together with the results of SD v2.1 depicted in Fig. 3, RID demonstrates consistent trends across all pre-trained diffusion models, with the defense performance improving as the perturbation scale increases. These results further validate RID’s generalizability and effectiveness across diverse model architectures.
@@
Analysis of perturbations generated by RID
Figure 4: Analysis of the perturbations generated by RID. The pixel values are normalized from $[0,255]$ to $[-1,1]$. Two Gaussian noise baselines with standard deviations of the perturbation level divided by $3$ (small, following the 3-sigma rule) and $1$ (large) are used for comparison, discretized to match the perturbation levels and normalized pixel range. a, Every perturbation is flattened as a vector to count the values of all elements. The distributions of RID perturbations differ markedly from those of Gaussian noise. b, Visualization of RID perturbations as images reveals structured, semantic patterns aligned with facial features, similar to those of optimization-based methods, while Gaussian noise appears random and unstructured. Perturbations for all methods are displayed at a scale of $1$ (e.g., multiplied by a factor of $255/8$) for clarity. c, RID perturbations consistently increase diffusion losses compared to Gaussian noise across various pre-trained (SD) and personalized (LoRA+TI and DB) diffusion models at all timesteps, suggesting the effectiveness of learned perturbations over natural noise. All facial images used in this figure are sourced from publicly available datasets and are permitted for academic purposes.
This section examines the properties of perturbations generated by RID in comparison with both Gaussian noise baselines and existing optimization-based methods. The corresponding results are reported in Fig. 4, where RID produces semantically meaningful perturbations that are significantly different from Gaussian noise and behave similarly to those produced by optimization-based methods, such as Anti-DB [19] and AdvDM [17].
@@
The pixel distributions of RID-generated perturbations, as shown in Fig. 4a, differ significantly from those of Gaussian noise baselines. All pixel values are normalized to [--1, 1], with two Gaussian baselines included for comparison: small Gaussian noise (with a standard deviation of one-third of the perturbation level, following the 3-sigma rule) and large Gaussian noise (matching the full perturbation level). RID perturbations exhibit a heavy-tailed statistical distribution, distinct from the symmetric distribution of Gaussian noise (even with truncation). This characteristic highlights RID’s ability to learn structured perturbations that deviate from the random and unstructured patterns seen in Gaussian noise.
@@
As shown in Fig. 4b, RID-generated perturbations reveal structured and semantic patterns that align closely with facial features, similar to those produced by optimization-based methods (e.g., Anti-DB and AdvDM). These perturbations focus on regions containing meaningful information, such as facial contours, to effectively disrupt identity replication. In contrast, Gaussian noise lacks any semantic structure, appearing random and uncorrelated with the content of the image. The semantic alignment of RID perturbations reflects the learned nature of its defense mechanism, ensuring that sensitive features are targeted for protection.